# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 5 Analysis of blockers and splitting
If $|I|\geq 3$, then we apply Lemma 21. So let $I=\{x,y\}$, where $\deg(x)\leq\deg(y)$ and let $A=N(x,y)$. Note that $|A|=\deg(x)+\deg(y)-4\geq\deg(x)-1$. The vertices in $A$ cannot form a clique, as then $x$ would have a funnel. So consider vertices $z_{1},z_{2}\in A$ with $z_{1}\not\sim z_{2}$.
If $z_{1}$ and $z_{2}$ are both subquartic, then we split on $I$; the subproblem $\langle G-N[I],k-|N[I]|-2$ has drop $(1,4)$ directly. The subproblem $G-I=G-\{x,y\}$ has non-adjacent subquadratic vertices $z_{1},z_{2}$, so by Observation 16, we have $R(x,y)\geq 2$ and subproblem $\langle G-I,k-|I|\rangle$ has drop $(1,4)$ after simplification.
So suppose $\deg(z_{1})\geq 5$. If $\operatorname{shad}(N[z_{1}])\geq 0$, then split on $z_{1}$; by Proposition 20, the subproblem $\langle G-z_{1},k-1\rangle$ has drop $(0.5,3)$, while the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(2,5)$ directly. Otherwise, suppose $\operatorname{shad}(N[z_{1}])\leq-1$ and $t$ is a blocker of $z_{1}$. Note that $t\notin I$ since $t\not\sim z_{1}$ whereas $z_{1}\in N(x)\cap N(y)$. We then apply (B) to $t,z_{1}$; the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(1,5)$ directly, and by Proposition 20, the subproblem $\langle G-\{t,z\},k-2\rangle$ has drop $(1,4)$. ∎
5.2 Branching rules for blocked vertices.
We now turn to developing branching rules for blocked vertices of various types. We need to work our way from low to high degree.
@@
Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 5-\deg(u)$ and $u$ has a blocker $x$ of degree at least 4. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,2),(2,5)]$.
@@
If $\operatorname{shad}(N[x])\leq 3-\deg(x)$ and $J$ is a min-set of $G-N[x]$, then $\operatorname{surp}_{G}(J\cup\{x\})\leq(3-\deg(x))+(\deg(x)-1)\leq 2$ and we can apply Lemma 22 to indset $J\cup\{x\}$.
So suppose $\operatorname{shad}(N[x])\geq 4-\deg(x)$. In this case, we apply (B) to $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2)$ directly. If $\deg(x)\geq 5$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1.5,5)$ directly. If $\deg(x)=4$, then consider a min-set $I$ of $G-N[u]$ with $x\in I$. We have $\operatorname{surp}_{G-N[x]}(I\cup\{u\}\setminus\{x\})\leq\operatorname{surp}_{G-N[u]}(I)+\deg(u)-\deg(x)\leq 1$. So $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1.5,5)$ after simplification. In either case, $\langle G-\{u,x\},k-2\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$ give drops $(1.2),(1.5,5)$ respectively. This branch-seq $[(1,2),(1.5,5)]$ dominates $[(0.5,2),(2,5)]$. ∎
@@
Case I: $\boldsymbol{\operatorname{shad}(N[u])\geq 5-\deg(u)}$. Splitting on $u$ gives subproblems $G-u,G-N[u]$ with drops $(0.5,1),(2,5)$ directly; the former has drop $(0.5,2)$ after simplification due to its 2-neighbor $z$.
@@
Case III: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{R(u,x)\geq 2}$. We apply (B) to vertices $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification, and by Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. We get branch-seq $[(1,4),(1,4)]$, which dominates $[(1,3),(1,5)]$.
Case IV: $\boldsymbol{G-N[u]}$ has some non-singleton min-set $\boldsymbol{I}$. Let $x\in I$. Necessarily $\deg(x)=3$, else it would be covered in Case II. We apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification due to subquadratic vertex $z$. By Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$ after simplification.
Case V: No previous cases apply. Suppose that no vertices in the graph are covered by any of the previous cases. Let $U$ denote the non-empty set of vertices $u$ with $\deg(u)\geq 5$ and $u$ having a 3-neighbor. We claim that every vertex $u\in U$ has degree exactly 5. For, suppose $\deg(u)\geq 6$, and since we are not in Case I we have $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq-2$. Then necessarily a min-set of $G-N[u]$ would have size at least two, and it would be covered in Case IV.
So every $u\in U$ has degree exactly 5, and is blocked by a singleton 3-vertex $x$, i.e. $N(x)\subseteq N(u)$. Let us say in this case that $x$ is linked to $u$; we denote by $X$ the set of all such 3-vertices linked to any vertex in $U$. We claim that each $x\in X$ has at least two neighbors of degree at least 5, which must be in $U$ due to their 3-neighbor $x$. For, suppose that $x$ has two subquartic neighbors $z_{1},z_{2}$; necessarily $z_{1}\not\sim z_{2}$ since $G$ is simplified and $x$ has degree 3. Then $R(u,x)\geq 2$ by Observation 16, which would have been covered in Case III.
On the other hand, we claim that each vertex $u\in U$ has at most one neighbor in $X$. For, suppose $u$ is linked to $x$ and $u$ has two 3-neighbors $x_{1},x_{2}\in X$, which must be non-adjacent since $G$ is simplified. Then Observation 16 gives $R(u,x)\geq 2$ due to subquadratic vertices $x_{1},x_{2}$. This would have been covered in Case III.
Thus, we see that each vertex in $X$ has at least two neighbors in $U$ and each vertex in $U$ has at most one neighbor in $X$. Hence $|U|\geq 2|X|$, and so by the pigeonhole principle, there must be some vertex $x\in X$ which is linked to two vertices $u_{1},u_{2}\in U$. We apply (B) to $u_{1},u_{2},x$, getting subproblems $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$. By Proposition 19, the latter subproblem has drop $(1,4)$. Furthermore, if $z_{1},z_{2}$ are two 5-neighbors of $x$, then $z_{1}\not\sim z_{2}$ else $x$ would have a 3-triangle. So $G-\{u_{1},u_{2},x\}$ has non-adjacent 2-vertices $z_{1},z_{2}$, and subproblem $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ has drop $(1,5)$ after simplification. ∎
@@
Suppose a vertex $u$ has $\deg(u)\geq 4$ and $\operatorname{shad}(N[u])\leq 4-\deg(u)$. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
@@
Let $I$ be a min-set of $G-N[u]$. By Observation 18, there is a vertex $x\in I$ which shares some neighbor $t$ with $u$. If $\deg(x)\geq 4$, then we apply Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. If $\deg(x)=3$ and $\deg(t)\geq 5$, we apply Lemma 24 to $t$. If $\deg(x)=3,\deg(t)\leq 4,|I|\geq 2$, we apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification (due to subquadratic vertex $t$) and by Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$.
So finally take $I=\{x\}$ where $\deg(x)=3$ and $x$ shares two neighbors $t_{1},t_{2}$ with $u$. Necessarily $t_{1}\not\sim t_{2}$ since $G$ has no 3-triangles. If either $t_{1}$ or $t_{2}$ has degree 5 or more, we can apply Lemma 24. If they are both subquartic, then we apply (B) to $u,x$; by Proposition 19, subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$ and by Observation 16 there holds $R(u,x)\geq 2$ and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification. Here, $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. ∎
@@
Suppose vertex $u$ has $\deg(u)\geq 5$ and $\operatorname{shad}(N[u])\leq 5-\deg(u)$.
@@
Let $I$ be a min-set of $G-N[u]$ of smallest size. If any vertex in $I$ has degree at least 4, the result follows from Proposition 23 (noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. So suppose that all vertices in $I$ have degree 3. By Observation 18, we may select some $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$. For this vertex $x$, define $Z=N(x)\cap N(u)$ and $Y=N_{G-N[u]}(x)$, where $|Z|+|Y|=\deg(x)=3$ and $|Z|\geq 1$.
@@
The vertices in $Z$ must be independent, as otherwise $G$ would have a 3-triangle with $x$. If $\operatorname{codeg}(z,z^{\prime})\geq 3$ for any vertices $z,z^{\prime}\in Z$, then $\operatorname{shad}(N[z])\leq 0$ (since $G-N[z]$ has a 1-vertex $z^{\prime}$), and we can apply Proposition 25. So we suppose the vertices in $Z$ share no neighbors besides $u$ and $x$. In this case, in the graph $G-\{u,x\}$, the vertices in $Z$ become subquadratic and share no common neighbors. By Proposition 12, we thus have $R(u,x)\geq|Z|$.
We next claim that if $|I|\geq 2$, then $R(u,x)\geq|Y|$. For, every vertex $y\in Y$ must have a neighbor $\sigma(y)\in I\setminus\{x\}$, as otherwise we would have $\operatorname{surp}_{G-N[u]}(I^{\prime})\leq\operatorname{surp}_{G-N[u]}(I)$ for non-empty ind-set $I^{\prime}=I\setminus\{x\}$, contradicting minimality of $I$. The vertices $\sigma(y):y\in Y$ must be distinct, as a common vertex $x^{\prime}=\sigma(y)=\sigma(y^{\prime})$ would have $\operatorname{codeg}(x,x^{\prime})\geq 2$ which we have already ruled out. So, in the graph $G-N[x]$, each vertex $\sigma(y)\in I$ loses neighbor $y$. By Observation 15, we have $\operatorname{shad}(N[x])\geq 3-\deg(x)=0$. So $G-N[x]$ has $|Y|\leq 2$ non-adjacent subquadratic vertices. Observation 16 implies that $R(N[x])\geq|Y|$.
At this point, our strategy is to apply (B) to $u,x$. If $|I|=1$, then either $\deg(u)=5,|Z|\geq 2$ or $\deg(u)\geq 6,|Z|=3$. Subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3)$ directly and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification; that is, it has drop $(1,4)$ if $\deg(u)=5$ and drop $(1,5)$ otherwise.
Otherwise, if $|I|\geq 2$, then subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification and subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3+|Y|)$ after simplification. Since $|Z|+|Y|=3$ and $|Z|\geq 1$, this always gives drops $(1,3),(1,5)$ or $(1,4),(1,4)$. ∎

### v7

@@ 3.1 Branching rules for indsets with surplus two
If $|I|\geq 3$, then we apply Lemma 21. So let $I=\{x_{1},x_{2}\}$ and let $A=N(x_{1})\cap N(x_{2})$. Note that $|A|=\deg(x_{1})+\deg(x_{2})-|N(I)|\geq\deg(x_{1})-1\geq 2$. The vertices in $A$ cannot form a clique, as then $x_{1}$ would have a funnel. So consider vertices $z_{1},z_{2}\in A$ with $z_{1}\not\sim z_{2}$.
If $z_{1}$ and $z_{2}$ are both subquartic, then we apply (B-Crit) to $I$; the subproblem $\langle G-N[I],k-4\rangle$ has drop $(1,4)$ directly. The subproblem $\langle G-I,k-2\rangle$ has non-adjacent subquadratic vertices $z_{1},z_{2}$, so by Observation 12 it has drop $(1,4)$ after simplification.
So suppose $\deg(z_{1})\geq 5$. If $\operatorname{shad}(N[z_{1}])\geq 0$, we split on $z_{1}$; by Observation 19, the subproblem $\langle G-z_{1},k-1\rangle$ has drop $(0.5,3)$ after simplification, while the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(2,5)$ directly. Otherwise, suppose $\operatorname{shad}(N[z_{1}])\leq-1$ and $t$ is a blocker of $z_{1}$. Note that $t\notin I$ since $t\not\sim z_{1}$ whereas $z_{1}\in N(x_{1})\cap N(x_{2})$. We then apply (B-Block) to $t,z_{1}$; the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(1,5)$ directly, and by Observation 19, the subproblem $\langle G-\{t,z\},k-2\rangle$ has drop $(1,4)$ after simplification. ∎
3.2 Branching rules for blocked vertices
We now turn to developing branching rules for blocked vertices of various types.
@@ 3.2 Branching rules for blocked vertices
Suppose $\operatorname{shad}(N[u])\leq 5-\deg(u)$ and $u$ has a blocker $x$ of degree at least 4. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,2),(2,5)]$.
@@
If $\operatorname{shad}(N[x])\leq 3-\deg(x)$ and $J$ is a min-set of GN[x]G-N[x], then $\operatorname{surp}_{G}(J\cup\{x\})\leq(3-\deg(x))+(\deg(x)-1)\leq 2$ and we can apply Lemma 22 to indset $J\cup\{x\}$.
So suppose $\operatorname{shad}(N[x])\geq 4-\deg(x)$. In this case, we apply (B-Block) to $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2)$ directly. If $\deg(x)\geq 5$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1.5,5)$ directly. If $\deg(x)=4$, then consider a min-set II of GN[u]G-N[u] with $x\in I$. We have $\operatorname{surp}_{G-N[x]}(I\cup\{u\}\setminus\{x\})\leq\operatorname{surp}_{G-N[u]}(I)+\deg(u)-\deg(x)\leq 1$. So $S(N[x])\geq 1$ and $G^{\prime}$ has drop $(1.5,5)$ after simplification. In either case, $\langle G-\{u,x\},k-2\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$ give drops $(1.2),(1.5,5)$ respectively. This branch-seq $[(1,2),(1.5,5)]$ dominates $[(0.5,2),(2,5)]$. ∎
@@
Case I: $\boldsymbol{\operatorname{shad}(N[u])\geq 5-\deg(u)}$. Splitting on $u$ gives subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$ with drops $(0.5,1),(2,5)$ directly; the former has drop $(0.5,2)$ after simplification due to its 2-vertex $z$.
@@
Case III: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{S(u,x)\geq 2}$. We apply (B-Block) to vertices $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification, and by Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. We get branch-seq $[(1,4),(1,4)]$, which dominates $[(1,3),(1,5)]$.
Case IV: $\boldsymbol{G-N[u]}$ has a non-singleton min-set $\boldsymbol{I}$. Let $x\in I$. Necessarily $\deg(x)=3$, else it would be covered in Case II. We apply (B-Block) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification due to subquadratic vertex $z$. By Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$ after simplification.
Case V: No previous cases apply. Suppose that no vertices in the graph are covered by any of the previous cases. Let $U\neq\emptyset$ denote the set of vertices $u$ with $\deg(u)\geq 5$ and $u$ having a 3-neighbor. We claim that every vertex $u\in U$ has degree exactly 5. For, suppose $\deg(u)\geq 6$, and since we are not in Case I we have $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq-2$. Then necessarily a min-set of GN[u]G-N[u] would have size at least two, and it would be covered in Case IV.
So every $u\in U$ has degree exactly 5, and is blocked by a singleton 3-vertex $x$, i.e. $N(x)\subseteq N(u)$. Let us say in this case that $x$ is linked to $u$; we denote by $X$ the set of all such 3-vertices linked to any vertex in $U$. We claim that each $x\in X$ has at least two neighbors of degree at least 5, which must be in $U$ due to their 3-neighbor $x$. For, suppose that $x$ has two subquartic neighbors $z_{1},z_{2}$; necessarily $z_{1}\not\sim z_{2}$ since $G$ is simplified and $x$ has degree 3. Then $S(u,x)\geq 2$ by Observation 12, which would have been covered in Case III.
On the other hand, we claim that each vertex $u\in U$ has at most one neighbor in $X$. For, suppose $u$ is linked to $x$ and $u$ has two 3-neighbors $x_{1},x_{2}\in X$, which must be non-adjacent since $G$ is simplified. Then Observation 12 gives $S(u,x)\geq 2$ due to subquadratic vertices $x_{1},x_{2}$. This would have been covered in Case III.
Thus, we see that each vertex in $X$ has at least two neighbors in $U$ and each vertex in $U$ has at most one neighbor in $X$. Hence $|U|\geq 2|X|$, and so by the pigeonhole principle, there must be some vertex $x\in X$ which is linked to two vertices $u_{1},u_{2}\in U$. We apply (B-Block) to $u_{1},u_{2},x$, getting subproblems $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$. By Proposition 20, the latter subproblem has drop $(1,4)$. Furthermore, if $z_{1},z_{2}$ are two 5-neighbors of $x$, then $z_{1}\not\sim z_{2}$ else $x$ would have a 3-triangle. So $G-\{u_{1},u_{2},x\}$ has non-adjacent 2-vertices $z_{1},z_{2}$, and subproblem $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ has drop $(1,5)$ after simplification. ∎
@@
Suppose a vertex $u$ has $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq 0$. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
@@
Let $I$ be a min-set of GN[u]G-N[u]. There must be some vertex $x\in I$ which shares some neighbor $t$ with $u$, as otherwise we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$, contradicting that $G$ is simplified. If $\deg(x)\geq 4$, we apply Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. If $\deg(x)=3$ and $\deg(t)\geq 5$, we apply Lemma 24 to $t$. If $\deg(x)=3,\deg(t)\leq 4,|I|\geq 2$, we apply (B-Block) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification (due to subquadratic vertex $t$) and by Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$.
So finally take $I=\{x\}$ where $\deg(x)=3$ and $x$ shares two neighbors $t_{1},t_{2}$ with $u$. Necessarily $t_{1}\not\sim t_{2}$ since $G$ has no 3-triangles. If either $t_{1}$ or $t_{2}$ has degree 5 or more, we can apply Lemma 24. If they are both subquartic, then we apply (B-Block) to $u,x$; by Proposition 20, subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$ and by Observation 12 the subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification. Here, $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. ∎
@@
Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 5-\deg(u)$.
@@
Let $I$ be an inclusion-wise-minimal min-set II of GN[u]G-N[u]. If any vertex in $I$ has degree at least 4, the result follows from Proposition 23 (noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. So suppose that all vertices in $I$ have degree 3.
There must be some vertex $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$ as otherwise we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$. For this vertex $x$, define $Z=N(x)\cap N(u)$ and $Y=N(x)\setminus N(u)$, where $|Z|+|Y|=\deg(x)=3$ and $|Z|\geq 1$.
@@
The vertices in $Z$ must be independent, as otherwise $G$ would have a 3-triangle with $x$. If $\operatorname{codeg}(z,z^{\prime})\geq 3$ for any vertices $z,z^{\prime}\in Z$, then $\operatorname{shad}(N[z])\leq 0$ (due to vertex $z^{\prime}$), and we can apply Proposition 25. So we suppose the vertices in $Z$ share no neighbors besides $u$ and $x$. In this case, in the graph $G-\{u,x\}$, the vertices in $Z$ have degree two and share no common neighbors. By Proposition 11, we thus have $S(u,x)\geq|Z|$.
Now suppose that $|I|=1$, i.e. $I$ consists of a single vertex $x$ with $\deg_{G-N[u]}(x)\leq 6-\deg(u)$. In particular, if $\deg(u)=5$, then $|Z|\geq 2$ and if $\deg(u)\geq 6$ then $|Z|\geq 3$. In this case, when we apply (B-Block) to $u,x$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3)$ directly and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ or $(1,5)$ after simplification, for $\deg(u)=5$ and $\deg(u)=6$ respectively. This shows the claimed bound.
So suppose that $|I|\geq 2$. Each vertex $y\in Y$ must have a neighbor $\sigma(y)\in I\setminus\{x\}$, as otherwise we would have $\operatorname{surp}_{G-N[u]}(I^{\prime})\leq\operatorname{surp}_{G-N[u]}(I)$ for non-empty indset $I^{\prime}=I\setminus\{x\}$, contradicting minimality of $I$. The vertices $\sigma(y):y\in Y$ must be distinct, as a common vertex $x^{\prime}=\sigma(y)=\sigma(y^{\prime})$ would have $\operatorname{codeg}(x,x^{\prime})\geq 2$ which we have already ruled out. So, in the graph GN[x]G-N[x], each vertex $\sigma(y)\in I$ loses neighbor $y$. By Observation 9, we have $\operatorname{shad}(N[x])\geq 3-\deg(x)=0$. So GN[x]G-N[x] has $|Y|\leq 2$ non-adjacent subquadratic vertices. Proposition 10 implies that $S(N[x])\geq|Y|$.
Finally, our strategy is to apply (B-Block) to $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification and subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3+|Y|)$ after simplification. Since $|Z|+|Y|=3$ and $|Z|\geq 1$, this always gives $(1,3),(1,5)$ or $(1,4),(1,4)$. ∎